OpenStudy (anonymous):

find all radian solutions in the interval 0<= x < 2pi round to four decimal places. 4sin^2x-12sinx-1=0 i did the quadratic formula and got 12 +- square root of 160, which breaks down into 12 +- 4 square root 10. im not sure how im supposed to get that into radians, or how ro round it 4 decimal places. am i doing this right?

OpenStudy (freckles):

\[4\sin^2(x)-12 \sin(x)-1=0 \\ \sin(x)=\frac{-(-12) \pm \sqrt{(-12)^2-4(4)(-1)}}{2(4)} \\ \sin(x)=\frac{12 \pm \sqrt{144+16}}{8} \\ \sin(x) =\frac{12 \pm \sqrt{160}} {8} \\ \sin(x)=\frac{12 \pm 4 \sqrt{10}}{8} \\ \sin(x)=\frac{3 \pm \sqrt{10}}{2} \] ok just wanted to check your answer so far now assume a is between -1 and 1 and sin(x)=a then \[x=\arcsin(a)+2n \pi \text{ or } x=-\arcsin(a)+(2n+1)\pi\] where n is an integer

OpenStudy (freckles):

however you should see (3+sqrt(10))/2 >1

OpenStudy (freckles):

so you only have to solve \[\sin(x)=\frac{3 - \sqrt{10}}{2}\]

OpenStudy (anonymous):

Thank you so much! That helped me out alot!

OpenStudy (freckles):

np